# AI Rules? Characterizing Reddit Community Policies Towards AI-Generated Content

### v1

@@ 3.2.2. Applying Subreddit Classification Labels
We next sought to label the subreddits in the Rules Subreddit Set with details about the type of community that each subreddit was home to. To this end, we devised a process to apply a Topic label using a taxonomy from Fiesler et al. (Fiesler et al., 2018) and Community Archetype labels using a taxonomy from Prinster et al. (Prinster et al., 2024). In applying our labels, we considered a subreddit’s name, topic, description, and rules metadata fields.
[…]
3.2.3. Identifying Emergent AI Rule Types
In order to better understand the types of rules that subreddits use to govern AI, we identified these rules in our sample and used an iterative inductive coding process to develop a descriptive taxonomy. To do this coding, we slightly modified the prompt developed and validated in Section 3.2.1 so that it would flag individual rules rather than a set of rules as being AI-related. Since the prompt required only a slight modification from the prior prompt, and since human coders would be manually reviewing the labels in the next step, we decided that we did not need to validate model performance before tagging the entire sample. We used this prompt to instruct gpt-4o-mini to label each of the $n=33,780$ individual rules in the AI Rules Subreddit Set, which produced the AI Rules Dataset ($n=4,150$) of individual rules about AI.
Next, two of the authors collaborated to identify emergent labels for the rules. The authors randomly selected $n=25$ rules from this dataset and independently performed affinity mapping (Harboe and Huang, 2015) by grouping rules that appeared similar into emergent categories. The authors then compared results and consolidated their categories into labels that best captured the elements of interest. Each of the two authors then independently applied this set of labels to another random sample of $n=25$ rules, once again grouping similar rules, including into new emergent groups where the existing labels did not suffice. The two authors repeated this process of randomly sampling, independently labeling, collaboratively comparing their results, and consolidating their labels, until no new labels emerged and they were confident that their taxonomy covered the relevant dimensions of variation in the dataset. The final label groupings were Rationale, Stance, Medium and Framing, which are listed alongside some of their possible labels in Table 2. We applied these labels to the entire AI Rules Dataset, as we describe next.
3.2.4. Applying Rule Classification Labels
We used LLMs to apply our emergent labels and an additional Rule Type label from Fiesler et al. (Fiesler et al., 2018) to the entire AI Rules Dataset. The two first authors collaboratively developed a codebook with instructions for applying these labels. Fiesler et al. does not provide codebook instructions, so the authors interpreted the examples presented in the study and came up with instructions for how to apply each of the $24$ possible labels 555“Rule Type” labels are not mutually exclusive, so we represented them as a series of binary labels, one for each possible Rule Type..
@@ 4. Findings
We group our findings by research question and present in sequence: (1) the prevalence of rules governing AI use, (2) the types of communities that have these rules, and (3) the types of rules that we find for governing AI use.
4.1. RQ1: How Common are Subreddit Rules Governing the Use of AI?
Table 3. Presence of Subreddit Community Rules Governing the Use of AI
N %
Rules Subreddit Set 80,290
w/ AI rules in 2023 crawl 1,411 1.76%
w/ AI rules in 2024 crawl 2,551 3.18%
YoY Increase 1,140 +81%
Top 1% Of Rules Subreddit Set by Subscriber Count 803
w/ AI rules in 2023 crawl 82 10.21%
w/ AI rules in 2024 crawl 152 18.93%
YoY Increase 70 +85%
@@ 4.1. RQ1: How Common are Subreddit Rules Governing the Use of AI?
Table 3 shows the results of our detection task described in Section 3.2.1. We find that during our first crawl in 2023, $n=1,411$ subreddits in the Rules Subreddit Set had rules governing the use of AI ($1.76\%$). By the time of our second crawl, this number increased by $81\%$ to $n=2,551$ ($3.18\%$). Rules governing the use of AI are more common in larger subreddits. For example, of the largest $1\%$ of subreddits in the Rules Subreddit Set by subscriber count, $10.21\%$ had AI rules during our first crawl. This percentage increased to $18.93\%$ in our second crawl. Figure 2 shows differences in the prevalence of AI rules for subreddits with different numbers of subscribers using data from the Rules Subreddit Sample. The figure contains two subplots which both have subscriber count split into deciles along the X axis (the labels represent the lower edge of the decile). The subplot on the left charts the percentage of subreddits with AI rules during our second crawl. The subplot on the right charts the percent increase in subreddits with AI rules between the two crawls. For example, the rightmost bar in the left subplot shows that in the largest decile of subreddits, about $11\%$ had rules about AI during the second crawl. The figure shows that subreddits with more subscribers are more likely to have AI rules, and that communities of all sizes saw increases in the prevalence of AI rules between our two crawls, though that increase was more prominent for larger communities.
@@
Figure 2. Percentage of subreddits with AI rules, bucketed by decile based on subscriber count. The first chart shows the percentage of subreddits with AI rules at the time of our second crawl and the second chart shows the percent change between the two crawls.
@@
The image contains two bar charts comparing the presence of AI-related rules in subreddits by subscriber count deciles:
Left Chart: ”Percentage of Subreddits With AI Rules”
This chart shows the percentage of subreddits with AI rules across different subscriber count deciles (measured by the lower edge of each decile). The subscriber counts (shown on the x-axis) are 11, 24, 48, 94, 188, 386, 850, 2155, 5978, and 22,345. The percentage of subreddits with AI rules steadily increases as subscriber count grows. The largest decile (22,345) has the highest percentage, above 10%, while the smallest decile (11) has a percentage near 0%. Right Chart: ”Increase in Subreddits With AI Rules”
This chart illustrates the year-over-year (YoY) percentage increase in the number of subreddits with AI rules, again grouped by subscriber count deciles. The subscriber counts (on the x-axis) are the same as the first chart: 11, 24, 48, 94, 188, 386, 850, 2155, 5978, and 22,345. The YoY increase is highest for the 94-subscriber decile, reaching almost 100%. Other significant increases are seen for the deciles of 188, 850, 5978, and 22,345, all of which have YoY increases around 80% to 90%.
@@ 4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
Table 4. Differences between subreddits in the Rules Subreddit Set ($n=80,290$) with and without AI rules. For numerical subreddit characteristics, we show the average value across the subreddits in each group. For binary characteristics, we show the percentage of subreddits with this characteristic in each group and denote these percentages with (%). While we computed differences for all 28 topics, we show here only the ten topics with the largest absolute differences.
@@
Metadata Subscriber count 33488.25 274350.19 240861.94***
Subreddit Age (Years) 5.91 8.06 2.15***
Subreddit created after ChatGPT (%) 5.88 9.88 4.0***
Community Archetype Content generation (%) 51.97 70.56 18.59***
Affiliation with an entity (%) 56.23 61.58 5.36***
Topical question and answer (%) 7.24 6.74 -0.5
Learning and perspective broadening (%) 17.9 14.11 -3.79***
Social support (%) 8.39 2.43 -5.96***
Topic Celebrity (%) 2.95 20.42 17.48***
Art (%) 2.59 13.84 11.25***
Video Games (%) 5.78 7.64 1.86***
Politics (%) 2.32 0.51 -1.81***
Business and Organizations (%) 2.86 0.67 -2.2***
Games (%) 8.82 6.12 -2.7***
Local (%) 4.03 1.25 -2.77***
Hobby (%) 7.93 4.59 -3.34***
Support (%) 7.55 1.76 -5.78***
Humor (%) 10.27 3.45 -6.82***
@@
To identify subreddit characteristics that predict the existence of AI rules, we perform a difference in means test between subreddits with and without AI rules. Table 4 presents the results of the test, showing various subreddit characteristics that differ between these two subreddit types. For numerical subreddit characteristics like Subscriber Count and Age the table shows the average value across these two groups of subreddits. For binary characteristics, such as whether a subeddit was created after the launch of ChatGPT, or if it has a specific Community Archetype or Topic label, we show the percentage of subreddits in each group with this characteristic. To test for significance we use chi-squared tests for the binary characteristics and the non-parametric Mann-Whitney U test for numerical characteristics. We implement Bonferroni correction to adjust for multiple comparisons.
@@
Table 4 exposes several key differences between subreddits with and without AI rules. Subreddits with AI rules tend to have more subscribers. These subreddits also tend to be newer and a higher percentage of subreddits with AI rules were created after ChatGPT was launched. We observe differences between the groups with respect to their association with the different Community Archetype labels. For example, there is a large and significant difference related to the Content Generation label; subreddits with this label are almost $20$ percentage points more likely to have AI rules. In contrast, subreddits of the Social Support Community Archetype are about $6$ percentage points less likely to have AI rules. Lastly, the Art and Celebrity topics are more common in subreddits with AI rules, encompassing about a third of all subreddits with AI rules compared to only $6\%$ of subreddits with no AI rules. These differences may indicate that communities focused on developing and sharing content are more likely to regulate AI usage in an attempt to maintain quality and standards, whereas communities that prioritize relationships among users are less concerned with the use of AI. We will further explore these implications in the discussion.
@@
We additionally run several logistic regression (LR) models to validate our results. As a baseline, we repeat the analysis done in Fiesler et al. (2018) and use a LR model to predict the presence of any set of rules for subeddits in the Broad Subreddit Set, using as input variables the following metadata gathered in our second crawl: Subscriber Count [^6] We log transform Subscriber Count to account for the variable’s skewed distribution., Subreddit Age (in years), and a binary indicator of whether or not the subreddit was created after the launch of ChatGPT. We compare these results with a LR model that specifically predicts the presence of AI rules, using as input all of the subreddit characteristics shown in Table 4. Comparing the two regression results isolates the subreddit characteristics that are specifically predictive of AI rules, not rules in general. We find that the Art and Celebrity Topic labels are both predictive of AI rules ($\beta=1.49,p<0.01$ and $\beta=2.12,p<0.01$), supporting the analysis above. In terms of Community Archetype labels, the Content Generation label is predictive of AI rules ($\beta=0.71,p<0.01$), while the Social Support label is negatively associated with the presence of AI rules ($\beta=-0.45,p<0.01$). These regression results are in line with the findings from our difference in means test.
@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
While our previous analyses have helped explain what types of communities are likely to explicitly regulate AI, we also consider differences in how communities regulate AI. Drawing on the AI Rules Set ($n=4,150$), we produce frequencies of the different Rule Type labels described in Sections 3.2.3 and 3.2.4. Table 5 shows the Rule Type labels used by Fiesler et al. (2018), while Table 6 presents the new labels that we produced via an iterative, inductive coding process. The tables show the percentage of rules in the set that were tagged with each label.
@@
Comparing our results to those of Fiesler et al. (Fiesler et al., 2018) illuminates some interesting differences between AI rules and general subreddit rules. Fiesler et al.’s results found that with general rules, Prescriptive rules, or rules that tell users what to do, occurred in about equal numbers as Restrictive rules that tell users what not to do [^7] These labels are not mutually exclusive: a rule could be tagged as both Prescriptive and Restrictive if it contains elements of both style.. However, in our AI Rules Set, $96.2\%$ of rules were tagged as Restrictive while only $79\%$ were tagged as Prescriptive. This further bears out in the data for our AI Rule: Stance label, which shows Unqualified Bans (e.g., “No AI-Generated Content”) as the most common label ($43.7\%$ of AI rules). Still, it is interesting to see that some communities are adopting stances other than outright bans, like the $32.2\%$ of AI rules labeled Stance: Qualified Ban and the $24.4\%$ labeled Stance: Disclosure. Perhaps this implies that communities will be flexible towards figuring out how to work with, rather than against, this new technology.
@@
Overall, we observe that a high number of AI rules are related to images. Table 5 shows that $50\%$ of rules were tagged with the Images label from Fiesler et al. (2018), which according to our codebook, indicates that the rule “is about posting visual content like images, photos, or paintings”. For comparison, Fiesler et al. (2018) tagged less than $8\%$ of the rules in their dataset with this label. We see slightly smaller numbers with our inductively coded labels, which used a codebook that looked for more specific language: Table 6 indicates that $36\%$ of AI rules are labeled as Medium: Images and $28\%$ are labeled Medium: Art. This observation further supports the finding in Section 4.2 that subreddits with Art and Celebrity topics are more likely to have AI rules, as these are usually image-based subreddits.
@@
Tables 5 and  6 also demonstrate that many AI rules on Reddit focus on concerns about content quality. Around $35\%$ of rules were labeled as relating to Low quality content; $17\%$ and $11\%$ respectively were labeled as Rationale: Low quality and Rationale: Low effort, meaning that they explicitly stated these concerns as their rationale for regulating AI use. These rationales are more prevalent than rationales based on accuracy, authenticity, and ethics. The focus on quality may suggest that explicit AI regulation is driven by practical concerns about content and reputation, not ideological objections over the use of AI more broadly. Notably, almost half of the AI rules were labeled Stance: Unqualified ban, indicating that many subreddits may find it easier to simply ban the use of AI outright, rather than trying to develop policies that can accommodate it under certain circumstances (e.g., requiring the disclosure of AI use).
@@
Table 5. Prevalence of Rule Type labels in the AI Rules Set ($n=4,150$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set.
@@
Content and behavior 99.78
Restrictive 96.19
Prescriptive 78.96
Images 50.48
Consequences, moderation, and enforcement 38.96
Low quality content 34.6
Post format 29.45
Copyright and piracy 17.23
Links and outside content 11.20
Spam 8.99
Off topic 7.3
NSFW 4.65
Reposting 4.34
Advertising and commercialization 3.08
Reddiquette and sitewide rules 1.01
Politics 0.94
Harassment 0.75
Doxxing and personal information 0.55
Spoilers 0.48
Hate speech 0.39
Trolling 0.39
Voting 0.1
Personal army 0.05
Personality 0.02
@@
Table 6. Prevalence of AI Rule Type labels in the AI Rules Set ($n=4,150$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set. Examples are drawn from the data but have been shortened for brevity.
AI Rule Type Percentage of AI Rules (%) Example
@@
Rationale Low quality 16.94 Low quality content will be removed. This includes AI generated images and other AI media.
Low effort 11.01 AI generated images or Chatgpt text content will be removed as “low effort” content. Please do not post AI generated content on this subreddit.
@@
Inauthentic 10.34 NO FAKES OR OVERLY SHOPPED PHOTOS - may result in a ban, serious offense. Images must be posted as published except for Quality AI-enhanced images. We ask that those titles include ”Enhanced.”
Original content 5.25 AI generated text does not qualify as an original source.
Off topic 3.4 Content generated by a computer algorithm is off topic and unwanted.
@@
Spam 3.13 Any form of spam / blogspam is not allowed. This includes links to book reviews, AI generated content, bots, social media channels and any affiliate links.
@@
Ethics 2.12 Given Miyazaki’s vehement opposition to AI and the ethical issues with AI generated art, posts that feature content generated by an AI model will be removed.
@@
Policy 1.47 Deep fakes violate reddit TOS. You will be banned if you post one.
Misleading 1.42 AI altered videos are misleading and will be removed.
@@
Inaccurate 0.96 AI is awesome, but when used to answer technical questions it often gives incorrect or sometimes dangerous recommendations.
@@
Inhuman 0.89 We believe Hip hop is human made, and for that reason we’ve decided to ban AI generated content.
Stance Unqualified ban 43.73 AI-generated content of any kind is forbidden.
@@
Qualified ban 32.24 AI Generated Fan Art is also not allowed, UNLESS it has been substantially edited by a human.
@@
Disclosure 24.36 If posting AI-generated art, specify in the title that it is AI-created, include what service used.
@@
Enabling 24.29 Recipes written by AI are ok but you must: 1) make the salsa, and 2) post a real photo of the completed salsa.
Requiring 0.92 Only post images generated by Ai, that’s it.
@@
Medium Content 35.76 On r/NFT, content that is fully generated by AI and not modified in any way shape or form is prohibited and will result in a ban from the community.
Images 35.73 Please no DALL·E mini or other AI generated images.
Art 27.93 No AI art allowed, as per our content policy.
@@
Fakes 21.88 No fake images of Leah Willamson are permitted, either photoshopped or AI generated
Unspecified 5.78 No AI - https://www.gofundme.com/f/protecting-artists-from-ai-technologies

### v2

@@ 3.3.4. Applying Subreddit Classification Labels
We also sought to label the subreddits in the Rules Subreddit Set with details about the type of community that each subreddit was home to. To this end, we devised a process to apply labels from two label categories used in prior HCI literature: Topic labels from Fiesler et al. (Fiesler et al., 2018) and Community Archetype labels from Prinster et al. (Prinster et al., 2024). In applying labels, we considered a subreddit’s name, description, and rules metadata fields.
@@ 4. Findings
We group our findings by research question and present in sequence: (1) changes in the prevalence of rules governing AI use, (2) the types of communities that have these rules, and (3) the types of rules that we find for governing AI use.
4.1. RQ1: How Has the Prevalence of Subreddit Rules Governing the Use of AI Changed Over Time?
Table 3. Prevalence of rules and rules governing the use of AI for subreddits in the Longitudinal Subreddit Set. We show the prevalence at the time of each crawl as well as the % change between the crawls. In addition to the prevalence in the entire dataset, we include the prevalence in the top 1% largest subreddits, ranked by subscriber count at the time of the 2024 crawl.
Jul 2023 Nov 2024
N % N % % Change
All subreddits in dataset ($N=227,737$)
had rules at time of crawl 66,505 29.2% 71,862 31.6% +8.1%
had AI rules at time of crawl 1,348 .6% 2,808 1.2% +108.3%
Top 1% largest subreddits ($N=2,278$)
had rules at time of crawl 2139 93.9% 2197 96.4% +2.7%
had AI rules at time of crawl 192 8.4% 390 17.1% +103.1%
@@ 4.1. RQ1: How Has the Prevalence of Subreddit Rules Governing the Use of AI Changed Over Time?
Table 3 shows the results of the detection task described in Section 3.3.1 on our Longitudinal Subreddit Set. We find that during the 2023 crawl, $N=1,348$ subreddits had rules governing the use of AI (.$6\%.$). By the time of our 2024 crawl, this number increased by $108.3\%$ to $N=2,808$ ($1.2\%$). Rules in general, and specifically rules governing the use of AI, are more common in larger subreddits. For example, $8.4\%$ of the largest $1\%$ of subreddits by subscriber count had AI rules during our first crawl. This percentage increased to $17.1\%$ in our second crawl. Figure 2 provides a more detailed demonstration of the differences in both the prevalence of AI rules and the increase in such rules among subreddits with different numbers of subscribers. The figure plots the number of subreddits in the Longitudinal Subreddit Set with AI rules at the time of each crawl. In order to expose the variation across subreddits of different sizes, we bucket subreddits along the X-axis into deciles based on their number of subscribers at the time of the second crawl (the labels represent the lower edge of the decile). Each decile contains approximately $22,773$ subreddits. For each decile we plot two bars representing the number of subreddits with AI rules: the left bar represents the number during the first crawl and the right bar represents the number during the second. Above the bars we include a numeric label representing the percent change between crawls. For example, the rightmost pair of bars shows that in the largest decile of subreddits, about $750$ had AI rules during the first crawl, and that this number increased by $128\%$ to about $1,750$ at the second crawl. The figure shows that subreddits with more subscribers are more likely to have AI rules, and that communities of all sizes saw increases in the prevalence of AI rules between our two crawls, though that increase was more prominent for larger communities.
For comparison, we additionally analyzed the prevalence of AI rules in subreddits in the Broad Subreddit Set that were created after our first crawl and were thus excluded from our Longitudinal Subreddit Set. The Broad Subreddit Set contained $N=30,962$ English language subreddits that were created after the first crawl (but were at least a month old during the second crawl). Of these newly created subreddits, $N=467$ had AI rules ($1.5\%$), higher than the percentage of subreddits with AI rules during the second crawl in the Longitudinal Subreddit Set ($1.2\%$). While this difference can be due to the fact that new subreddits have different characteristics, our separate analysis in the following section also suggests that newer subreddits are more likely to have AI rules.
@@
Figure 2. Number of subreddits in the Longitudinal Subreddit Set with AI rules during each of our crawls. To demonstrate differences across subreddit size, subreddits are bucketed into deciles based on their subscriber count at the time of the second crawl. Each decile contains approximately $22,773$ subreddits. The bars shows the number of subreddits with AI rules in each decile and the labels above the bars indicate the percent change between crawls.
@@
The image contains a bar chart showing the number of AI-related rules in subreddits during two different crawls, with subreddits bucketed into deciles by subscriber count. The Y-axis represents the number of subeddits and the X-axis is subscriber count decile at the time of the second crawl (measured by the lower edge of each decile). Each decile has two bars above it: the left bar represents the 2023 crawl and the right represents the 2024 crawl. The number of subreddits with AI rules steadily increases as subscriber count grows. The largest decile (7,241) has the highest number, above 1750 during the secon crawl, while the smallest decile (10) has a number near 0.
@@ 4.2. RQ2: What Types of Communities Have Rules Governing AI Use?
Table 4. Differences between subreddits in the Rules Subreddit Set ($N=99,969$) with and without AI rules. For numerical subreddit characteristics, we show the average value across the subreddits in each group. For binary characteristics, we show the percentage of subreddits with this characteristic in each group and denote these percentages with (%). While we computed differences for all 28 topics, we show here only the ten topics with the largest absolute differences.
@@
Metadata Subscriber count 37,865.69 27,3571.19 235705.5***
Subreddit Age (Years) 4.94 6.32 1.38***
Subreddit created after ChatGPT (%) 24.78 31.33 6.56***
Community Archetype Content generation (%) 51.68 73.75 22.07***
Affiliation with an entity (%) 57.47 63.3 5.83***
Topical question and answer (%) 7.29 6.52 -0.78
Learning and perspective broadening (%) 20.52 16.61 -3.91***
Social support (%) 10.29 4.38 -5.91***
Topic Celebrity (%) 3.87 18.33 14.45***
Art (%) 2.36 12.99 10.63***
Entertainment (%) 19.01 22.96 3.95***
Politics (%) 2.21 0.54 -1.67***
Business and Organizations (%) 2.71 0.99 -1.73***
Games (%) 6.97 4.68 -2.29***
Local (%) 4.22 1.11 -3.11***
Hobby (%) 8.07 4.7 -3.36***
Humor (%) 9.75 3.32 -6.43***
Support (%) 8.98 2.52 -6.46***
@@
To identify subreddit characteristics that suggest the existence of AI rules, we first perform a difference-in-means test between subreddits with and without AI rules. Table 4 presents the results of the test, showing various subreddit characteristics that differ between these two subreddit types. For numerical subreddit characteristics like Subscriber Count and Age the table shows the average value across these two groups of subreddits. For binary characteristics, such as whether a subeddit was created after the launch of ChatGPT or if it has a specific Community Archetype or Topic label, we show the percentage of subreddits in each group with this characteristic. To test for significance we use chi-squared tests for the binary characteristics and the non-parametric Mann-Whitney U test for numerical characteristics. We implement Bonferroni correction to adjust for multiple comparisons.
@@
Table 4 exposes several key differences between subreddits with and without AI rules. Subreddits with AI rules tend to have more subscribers. These subreddits also tend to be newer; correspondingly, a higher percentage of subreddits with AI rules were created after ChatGPT was launched. We observe differences between the groups with respect to their association with the different Community Archetype labels. For example, there is a large and significant difference related to the Content Generation label; subreddits with this label are about $22$ percentage points more likely to have AI rules. In contrast, subreddits of the Social Support Community Archetype are about $6$ percentage points less likely to have AI rules. Lastly, the Art and Celebrity topics are more common in subreddits with AI rules, encompassing about a third of all subreddits with AI rules compared to only $6\%$ of subreddits with no AI rules.
@@
We run several logistic regression (LR) models to validate our results. As a baseline, we repeat the analysis done in Fiesler et al. (2018) and use a LR model to predict the presence of any set of rules for subeddits in the Broad Subreddit Set, using as input variables the following metadata gathered in our second crawl: Subscriber Count [^6] We log transform Subscriber Count to account for the variable’s skewed distribution., Subreddit Age (in years), and a binary indicator of whether or not the subreddit was created after the launch of ChatGPT. We compare these results with a LR model that specifically predicts the presence of AI rules, using as input all of the subreddit characteristics shown in Table 4. Comparing the two regression results isolates the subreddit characteristics that are specifically predictive of AI rules, not rules in general. We find that the Art and Celebrity Topic labels are both predictive of AI rules ($\beta=1.47,p<0.01$ and $\beta=1.41,p<0.01$), supporting the analysis above. In terms of Community Archetype labels, the Content Generation label is predictive of AI rules ($0.010.83,p<$ 0.01). These regression results are in line with the findings from our difference-in-means test.
One possible explanation for why Content Generation communities are more likely to enact rules about AI could be because this type of community is generally less personal, with looser ties between members, and thus may need to rely on formal rules to maintain content quality and standards. Communities that prioritize relationships among users, on the other hand, may be able to use more personal methods to shape user behavior, such as informal norms. For example, consider a Wes Anderson subreddit whose purpose is to discuss “photos that accidentally share Wes Anderson’s unique style.” We labeled this subreddit with both the Art topic and the Content Generation community archetype. The language of this subreddit’s AI rules reflects a distinct concern with the veracity of posted images, rather than concerns about authentic interactions. The subreddit’s rules state that “Your photos don’t have to be images taken by you, but they must be real photos.” Similarly, subreddits labeled with the Celebrity topic tend to group AI rules with bans of photoshop, white borders, or low-resolution images, all of which are concerned with the format and quality of the subreddit’s content. Some Celebrity subreddits seem to encourage AI, but only in cases where it may produce enhanced images. In contrast, subreddits that we labeled with the Social Support community archetype like r/introverts – a subreddit that does not have an AI rule – emphasize in their rules that they care about “authentic connections.” We further explore the implications of these observations in Section 5.
@@ 4.3. RQ3: What Types of Rules Exist for Governing AI?
While our previous analyses explored what types of communities are likely to explicitly regulate AI, we also consider differences in how communities regulate AI. Drawing on the AI Rules Set ($N=4,458$), we produce frequencies of the different Rule Type labels described in Sections 3.3.2 and 3.3.3. Table 5 shows the Rule Type labels used by Fiesler et al. (2018), while Table 6 presents the new labels that we produced via an iterative, inductive coding process. The tables show the percentage of rules in the set that were tagged with each label.
@@
Comparing our results to those of Fiesler et al. (Fiesler et al., 2018) illuminates some interesting differences between AI rules and general subreddit rules. Fiesler et al.’s results found that with general rules, Prescriptive rules, or rules that tell users what to do, occurred in about equal numbers as Restrictive rules that tell users what not to do [^7] These labels are not mutually exclusive: a rule could be tagged as both Prescriptive and Restrictive if it contains elements of both style.. However, in our AI Rules Set, $96.84\%$ of rules were tagged as Restrictive while only $67.7\%$ were tagged as Prescriptive. A Restrictive rule might include statements like “No AI (chatGPT etc.)” or “No AI generated content” while a Prescriptive rule might instead tell users whether to cite if AI is used. This pattern further bears out in the data for our AI Rule: Stance label category, which shows Unqualified Bans (e.g., “No AI-Generated Content”) as the most common label ($55.23\%$ of AI rules). It seems that many subreddits may find it easier to simply ban the use of AI outright, rather than trying to develop policies that can accommodate it under certain circumstances (e.g., requiring the disclosure of AI use). Still, we do see evidence of some communities adopting stances other than outright bans, like the $24.23\%$ of AI rules labeled Stance: Qualified Ban and the $17.95\%$ labeled Stance: Disclosure. Perhaps this implies that some communities will be flexible towards figuring out how to work with, rather than against, this new technology.
@@
Overall, we observe that a high number of AI rules are related to images. Table 5 shows that $46.55\%$ of rules were tagged with the Images label from Fiesler et al. (2018), which according to our codebook, indicates that the rule “is about posting visual content like images, photos, or paintings”. For comparison, Fiesler et al. (2018) tagged less than $8\%$ of the rules in their dataset with this label. We see slightly smaller numbers with our inductively coded labels, which used a codebook that looked for more specific language: Table 6 indicates that $31.4\%$ of AI rules are labeled as Medium: Images and $30.26\%$ are labeled Medium: Art. This observation further supports the finding in Section 4.2 that subreddits with Art and Celebrity topics are more likely to have AI rules, as these are usually image-based subreddits.
@@
Tables 5 and  6 also demonstrate that many AI rules on Reddit focus on concerns about content quality. Around $28\%$ of rules were labeled as relating to Low quality content; $13\%$ and $11\%$ respectively were labeled as Rationale: Low quality and Rationale: Low effort, meaning that they explicitly stated these concerns as their rationale for regulating AI use. Rules about Low quality content emphasize the importance of “creative merit” and content that takes more than “only a few minutes of effort.” These rationales are more prevalent than rationales based on spam, ethics, or accuracy. The focus on quality may suggest that explicit AI regulation is driven by practical concerns about content and reputation, not ideological objections over the use of AI more broadly. This observation echoes the “content quality concerns” shared by moderators in Lloyd et al. (2023).
@@
Table 5. Prevalence of Fiesler et al.’s Rule Type labels in the AI Rules Set ($N=4,458$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules in this set.
@@
Content and behavior 99.73
Restrictive 96.84
Prescriptive 67.7
Images 46.55
Consequences, moderation, and enforcement 32.48
Low quality content 27.77
Post format 19.25
Copyright and piracy 16.49
Links and outside content 10.54
Spam 9.02
Off topic 7.11
Reposting 5.88
NSFW 5.59
Advertising and commercialization 3.93
Reddiquette and sitewide rules 1.05
Politics 1.03
Harassment 0.96
Spoilers 0.7
Hate speech 0.61
Doxxing and personal information 0.54
Trolling 0.43
Voting 0.18
Personal army 0.13
Personality 0.04
@@
Table 6. Prevalence of AI Rule Type labels in the AI Rules Set ($N=4,458$). We show the percentage of rules in the set that were tagged with each label. Rules may be tagged with multiple labels and each subreddit may have multiple rules included in this set. Examples are drawn from the data but have been shortened for brevity. See the tagging prompt in Appendix A.2 for label definitions.
Rule Category Rule Type % Of Rules Example
@@
Rationale Low quality 13.19 Low quality content will be removed. This includes AI generated images and other AI media.
@@
Inauthentic 11.93 NO FAKES OR OVERLY SHOPPED PHOTOS - may result in a ban, serious offense. Images must be posted as published except for Quality AI-enhanced images. We ask that those titles include ”Enhanced.”
Low effort 11.06 AI generated images or Chatgpt text content will be removed as ”low effort” content. Please do not post AI generated content on this subreddit.
Original content 7.0 AI generated text does not qualify as an original source.
Off topic 3.1 Content generated by a computer algorithm is off topic and unwanted.
@@
Spam 3.25 Any form of spam / blogspam is not allowed. This includes links to book reviews, AI generated content, bots, social media channels and any affiliate links.
@@
Ethics 2.98 Given Miyazaki’s vehement opposition to AI and the ethical issues with AI generated art, posts that feature content generated by an AI model will be removed.
@@
Policy 1.37 Deep fakes violate reddit TOS. You will be banned if you post one.
Misleading 1.7 AI altered videos are misleading and will be removed.
@@
Inaccurate 0.83 AI is awesome, but when used to answer technical questions it often gives incorrect or sometimes dangerous recommendations.
@@
Inhuman 0.9 We believe Hip hop is human made, and for that reason we’ve decided to ban AI generated content.
Stance Unqualified ban 55.23 AI-generated content of any kind is forbidden.
@@
Qualified ban 24.23 AI Generated Fan Art is also not allowed, UNLESS it has been substantially edited by a human.
@@
Disclosure 17.95 If posting AI-generated art, specify in the title that it is AI-created, include what service used.
@@
Enabling 18.51 Recipes written by AI are ok but you must: 1) make the salsa, and 2) post a real photo of the completed salsa.
Requiring 1.1 Only post images generated by Ai, that’s it
@@
Medium Content 52.74 On r/NFT, content that is fully generated by AI and not modified in any way shape or form is prohibited and will result in a ban from the community.
Images 31.4 Please no DALL·E mini or other AI generated images.
Art 30.26 No AI art allowed, as per our content policy.
@@
Fakes 15.79 No fake images of Leah Willamson are permitted, either photoshopped or AI generated
Unspecified 5.36 No AI - https://www.gofundme.com/f/protecting-artists-from-ai-technologies